Joe348:

i hate math 6(n+5)=2

brokenbaexd:

so ima help u @joe348 if u dont mind

Joe348:

cool

mercades74605:

dont you gotta like distribute

Joe348:

mhm

brokenbaexd:

yes

@mercades74605 wrote:
dont you gotta like distribute

brokenbaexd:

Joe u have to multiply .....6(n+5)

brokenbaexd:

do it and tell me what u get afterwards

Joe348:

woah woah woah

Joe348:

6n+30

brokenbaexd:

okay... good

@joe348 wrote:
6n+30

brokenbaexd:

then u subtract 30 from both sides of the equation

Joe348:

mhm

Joe348:

uh 6n=-28 right?

Joe348:

wait do i divide 6 next joshhhh

mercades74605:

yea i think so

Joe348:

-4.26?

Joe348:

wait i meant -4.67

brokenbaexd:

wait .....wait....wait .....

@joe348 wrote:
uh 6n=-28 right?
Joe thats right but after this u simply

Joe348:

nu nu nu you divide

Extrinix:

\(6(n+5)=2\) Distribute the 6. \(6(n+5)=2\) \(6n+30=2\) Subtract 30 from each side to try and get \(n\) by itself. \(6n+30=2\) \(6n=-28\) Solve it by dividing 6 on each side. \(6n=-28\) \(n=4.\overline{6}\) Now depending on how it wants you to answer, it could ask for the fraction. \(n=-4\dfrac{2}{3}\) OR \(n=-\dfrac{10}{3}\) OVERVIEW: - Distribute. - Subtract from both sides. - Divide on both sides to get \(n\) by itself.

mercades74605:

what type of math is this, algebra1?

Joe348:

idek

brokenbaexd:

yes i think so

@mercades74605 wrote:
what type of math is this, algebra1?

mercades74605:

not suppose to give direct answer

Extrinix:

It is Algebra I, yes.

Rylee88:

@Extrinix, You should know not to give out direct answers. You are level blue you know better. Don't give out direct answers.

Joe348:

exactly i was right.. tryna make me feel dumb .-.

mercades74605:

@extrinix wrote:
@rylee88 wrote:
@Extrinix, You should know not to give out direct answers. You are level blue you know better. Don't give out direct answers.
I didn't, I explained it, clearly.
and who bouta take time and read all that

Rylee88:

Still you could of put it piece by piece instead of that.

Joe348:

oh josh thxs you